OpenStudy (anonymous):

What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot? 0.25pi sq. ft. 0.5pi sq. ft. pi sq. ft

OpenStudy (imstuck):

area of a sector formula is this:

OpenStudy (anonymous):

Given a circle with radius r = 8 and a sector with subtended angle measuring 45°, find the area of the sector and the arc length. They've given me the radius and the central angle, so I can just plug into the formulas. For convenience, I'll first convert "45°" to the corresponding radian value of π/4: A = ((pi/4)/2)(8^2) = (pi/8)(8^2) = 8pi, s = (pi/4)(8) = 2pi area A = 8π, arc-length s = 2π Given a sector with radius r = 3 and a corresponding arc length of 5π, find the area of the sector. Copyright © Elizabeth Stapel 2010-2011 All Rights Reserved For this exercise, they've given me the radius and arc length. From this, I can work backwards to find the subtended angle. Then I can plug-n-chug to find the sector area. 5pi = (theta)(3), (theta) = (5/3)pi So the central angle is (5/3)π. Then the area of the sector is: A = ((5/3)pi / 2)*(3^2) = ((5/6)pi)*(9) = (15/2)pi A = (15 pi) / 2

OpenStudy (imstuck):



OpenStudy (anonymous):

thats an example

OpenStudy (anonymous):

Stuck do i have the right idea?

OpenStudy (imstuck):